Write down the relations among  the roots of the following equation a0x^3+a1x^2+a2x+a3=0

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
Let roots be b, c and d.
a+b+c= -a1/a0
ab+bc+ca= a2/a0
abc= -a3/a0
These are also known as the vietas formula